OpenStudy (anonymous):

jack mows the lawn in 4 hours. Marylin can mow the same lawn in 3 hours. How long would it take them to mow together

OpenStudy (anonymous):

I think its 1 5/7

OpenStudy (anonymous):

you're right

OpenStudy (anonymous):

\[\frac{t}{3} +\frac{t}{4} = 1\] \[4t + 3t = 12\] \[7t = 12\] \[t = \frac{12}{7}\] \[t = 1 \frac{5}{7}\] good job!

OpenStudy (anonymous):